OpenStudy (anonymous):

PLease help me will medal Which of the following expressions are equivalent to 2/(x^6-y^6) choose all that apply A 1/(x^3-y^3)*1/(x^3+y^3) B 2/(x^3-y^3)*1/(x^3+y^3) C 2/(x^3-y^3)*1/(x^3-y^3) D 2/(x^3)^2-(y^3)^2

OpenStudy (anonymous):

a,c,d

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

dca

OpenStudy (mathstudent55):

\(a^2 - b^2 = (a + b)(a - b) \) Now think of \(x^6 - y^6\) as \((x^3)^2 - (y^3)^2\) and factor using the factoring of a difference of squares shown above.

OpenStudy (mathstudent55):

\(x^6 - y^6 = (x^3)^2 - (y^3)^2 = (x^3 + y^3)(x^3 - y^3) \)

OpenStudy (mathstudent55):

\(\dfrac{2}{x^6 - y^6} = \dfrac{2}{(x^3 + y^3)(x^3 - y^3)} \)

OpenStudy (mathstudent55):

Among the choices, there are 2 correct answers.

OpenStudy (anonymous):

are they b and d @mathstudent55

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

it was b and d

OpenStudy (mathstudent55):

Yes, the correct answers are b and d.